OpenStudy (anonymous):

The ratio of white socks in a drawer to blue socks is 3:4. If there are 6 white socks in the drawer, how many blue socks are there? 8 7 6 4 3

OpenStudy (anonymous):

3/4 = 6/x Cross multiply: 24=3x x=8

OpenStudy (anonymous):

A lab has 65 mice. Some are white and the rest are gray. Which of the following could be the ratio of white mice to gray mice in the lab? 1 : 3 5 : 8 5 : 11 3 : 4 5 : 12